Let cos beta =a .Find the expression for cos 2beta and sin 2beta in terms of a
and hence confirm that - cos(square)2beta + sin(square)2beta = 1

Answer by stanbon(75887)

Let cos(beta) =a/1; this implies that x = a and r=1
Therefore y = sqrt(1-a^2)
Therefore sin(beta)= [sqrt(1-a^2)]/r
-------------------------------------
Find the expression for cos(2beta) and sin(2beta) in terms of a:
cos(2beta) = cos^2(beta)-sin^2(beta
= a^2 - [(1-a^2)/r^2]
------------------------------------ ------
sin(2beta) = 2*sin(beta)*cos(beta)
= 2 * sqrt(1-a^2)/r]
--------------------------
and hence confirm that - cos^2(2beta) + sin^2(2beta) = 1
------------------
Subsitute to confirm that statement is true.
====================
Cheers,
Stan H.